OpenStudy (anonymous):

simplify the expression: (8- squareroot of -1) - (-3 + squareroot of -16)

OpenStudy (mertsj):

8-i-(-3+4i)= 8-i+3-4i = 11-3i

OpenStudy (anonymous):

thank you. can you help me with two more possibly?

OpenStudy (mertsj):

Possibly

OpenStudy (anonymous):

-5(1+2i)+3i(3-4i)....isnt this -5+-10i + 6i-12i and then -5-16i right?

OpenStudy (mertsj):

-5-10i+9i-12i^2=-5-10i+9i+12=17-i

OpenStudy (anonymous):

okay i see where i messed up. thank you. and one more?

OpenStudy (mertsj):

ok

OpenStudy (anonymous):

(-2+ squareroot of -9)(6+ squareroot of -25) sorry I don't understand these.

OpenStudy (mertsj):